OpenStudy (anonymous):

Can anyone help me out on these two word problems? 17. ) I am thinking of two integers. One is 7 less than a number, and the other is 2 greater than the same number. a) write expressions for the two numbers. b) write a polynomial expression for the product of the numbers. 18.) Monica's current office is square. Her office in the company's new building will be 3 feet wider and 5 feet longer. a) write an expression for the dimensions of Monica's new office. b) write a polynomial expression for the area of Monica's new office. c) Suppose Monica's current office is 7 feet by 7 feet. How

OpenStudy (anonymous):

How much larger will her new office be?*

OpenStudy (lgbasallote):

for #17) let x be the number one is 7 less than x so one number would be x - 7 the other is 2 greater than x so the other number would be x+2 so the two integers are x - 7 and x + 2 does that help?

OpenStudy (lgbasallote):

for #18) let l be the original length of her office and w is the width of her original office the new office is 3 feet wider so w + 3 the new office is also 5 feet longer so l + 5 so the dimensions of her office will be w+ 3 by l + 5 does that help?

OpenStudy (anonymous):

idk if this is right but for the first one i got the two equations x-7=y y+2=x if you use substitution for these systems you get 2(x-7) = x then do that and you get x=14 plug it in to get y and y=7 so yeah there you go idk how you put that in the so called polynomial expression

OpenStudy (anonymous):

Thanks Ya'll.